A dishonest milkman sells milk at cost price but he mixes water and earns 50/3% profit. Find the ratio of milk and mixture in the mixture.  Please explain it logically.  Thank u.

Grade:8

2 Answers

Riddhish Bhalodia
8 years ago
say the milk costs C rupees per liter, and the dishonest milkman mixes x liters milk and 1-x liters water to make the one liter milk
He sells the mixed liter of milk at C rupees but the actual cost to him (assuming water doesn’t cost anything is xC rupees only)
Hence he makes a profit of
C-xC rupees = 50/3% of the cost which is xC
thus
(1-x)C = 50/300 \times xC
x = 6/7
hence milk:water = 6:1
Irfan khan
5 years ago
Profit is only from water and profit is 50/3 %. Water is 100% so the ratio of milk and water is-
 
100:50/3= 6:1 answer